UT 20161124 LUCI1 AO Nighttime Commissioning/FLAO-tests

Logs from JPower

Observer LBTO: JPower, JHill
Observer AO: DMiller
Telescope Operator: DHuerta
Mountain support: JMorris

Plan for the night

- AO/LUCI Comm when conditions suitable,

- LN for 1.5 hours at the start of the night

- Pointing model/D-time science other half if conditions not suitable for AO/LUCI Comm.

Summary

LN handed over at 18:40. The seeing conditions were marginal for AO. We performed LUCI AO
comm all night, taking shot LUCI images in various binning modes. Seeing was poor at points at
the start and a pointing model with LUCI LUCI was performed until the RFBG was removed from use
for the night. Completed Q0104 - Bin3 66 modes in H, Trapezium in Bin1 400 modes switching to 300 modes.
Investigations into NCPA analysis done and binning comparison. Moved on to 4C 41.17 bin 4 36 modes.
Attempted strehl measurement obs at the end of the night. Light Cirrus wafting through most of the night
getting thicker fro the last hour. Seeing ranged from 0.5-1.5 on the Dimm, fluctuating between 0.6\x94-1.0\x94 most of the night.

Issues

Temperature problem with Drive 1 motor temperature RFBG. Prevented some DX AO work and completion
of pointing model in LUCI LUCI. Suspect problems with NCPA analysis package. LUCI1 N30 Field stop shifted
between start of last night and tonight, Dave emailed about it.

Details

All times are UT
Files are luci1.20161024.NNNN.fits

23:20 LUCI1 darks
10 x (1 x 2.6) NORMAL #0001-0010
10 x (6 x 10) INTEGRATED #0011-0020
10 x (20 x 2.6) INTEGRATED #0021-0030
10 x (24 x 2.6) INTEGRATED #0031-0040

01:30 Seeing improved to 0.9\x94 Doug has been called to come in. 45 min out

01:41 Telescope handed over to D-time work.

03:26 AO preset sent

03:40 SX Communication WFS to Adsec issues (similar to those that were happening with DX2
earlier in the week). Also configuration issue with DX - Doug is progressing with that.

03:55 Seeing DIMM 0.93\x94 Guider 0.67\x94

04:05 Seeing degraded, highly variable. Switching back to pointing model. Going to pick up
where we left off

05:25 Authorized for LUCI1 only, going to try some LUCI AO scripts. DIMM~0.95\x94
Seeing is suddenly taking a turn for the better.

05:28 Sending preset to Q0104 - Bin3 66 modes 200 Hz
#0043 Ks Note that Full dither pattern was completed for this target in Ks
yesterday. Will skipped to H. #0044-0077 DIMM 0.78\x94, Guider 0.55\x94

06:28 #0078 loop opened when seeing blew up suddenly to 1.6\x94 on DIMM 1.04 on guider

06:32 Resending Preset in Ks #0079-0081 (left Ks in place).

06:45 H #0082-0106 picking up at the same position as loop opened previously.
DIMM 0.92\x94 Guider 0.69\x94 varying up to 0.91\x94 on guider 1.07\x94 on Dimm

07:22 Preset to Trapezium sent. Failed with GCS Error. David is doing pointing check.

07:27 Preset to Trapezium resent. Blind Blind #0107 Bin 1, 400 modes, 990Hz
GUIDER: 0.45\x94 DIMM: 0.6\x94. #0108 Open Loop failed.

07:40 Often failure was coming back from first offset with blind blind filter with absolute (0,40)
offset. ​Doug did some testing yesterday and found that with offsets on the order 30+" in Bin1
only, the loop closes after the offset then opens because of pupil alignment issues. This issue
is being investigated. I opted to remove the 40" offset from the Trapezium, ​Changing the offset
to absolute (0,0). The part of the script that takes Images with light on them (not blind blind
filters) were not altered.

07:32 Preset resent for Trapezium with updated script. #0109
#0110-0129 Bin 1, 400 modes, 990Hz NCPA On Guider 0.47\x94 DIMM 0.67\x94
#0125 Open Loop but continued without failure in script

08:19 Pausing at end of BrGam. Doug is making adjustments, optical gain not looking correct.

08:30 Redoing last map-point in Br Gamma but with Optical Gain is off, no NCPA
#0130-0131, 20161124.0131 Offset failed

08:39 Resending Trapezium preset #0132-0137

08:54 BrGam #0138-0142 Optical Gain is off, no NCPA Bin 1, 400 modes.
DIMM 0.81\x94 Guider 0.53\x94

09:01 NCPA back on. Continuing to next filter H2. NCPA, Bin 1, 400 modes.
#0143-0152, #0153 loop opened

09:16 Resend Preset to trapezium. NCPA, Bin 1, 300 modes. Seeing DIMM 0.68\x94 GUIDER 0.45\x94

09:28 H2 & FeII NCPA Bin1 300 modes #0154-0185

10:16 Preset sent to 4C 41.17. Bin 4, 36 modes freq 100 Hz
Ks #0186-0249 Guider 0.8\x94 DIMM 1.01\x94. Thicker cirrus is starting to move in.

12:07 Moving on to AO star for Strehl measurement obs.

12:26 Cirrus is thick enough that the star has about 2 mags of extinction from cirrus and has
been bumped to bin2. Need to find another option.

12:30 Moved to AO302 6.8 in R, 5.1 in H. Looks to be ~ 9.3 currently. 0.53\x94 guider 0.95\x94 on DIMM

#0257 copy issue
#0258 1 (50x0.16) 2\x94 offset background image
#0259 1 (50x0.16) Target Bin1 300 modes NCPA On
#0260 1 (50x0.16) Target Bin1 300 modes NCPA On
#0261 1 (50x0.16) Target Bin1 300 modes NCPA On
#0262 1 (50x0.16) 2\x94 offset background image
Resend preset
#0264 1 (50x0.16) Target Bin1 400 modes NCPA Off
#0265 1 (50x0.16) Target Bin1 400 modes NCPA Off
#0266 1 (50x0.16) Target Bin1 400 modes NCPA On *

13:16 Star cotton balled, seeing blew up. Seeing on guider 0.82\x94 No DIMM measurement
available. Optical gain issues.

13:25 Closing up. Cirrus too thick and seeing degraded too much for anything useful.

13:51 Luci internal imaging flats

N30 H2 #0277-0281 5 x (3 x 4.5)lamp off
N30 H2 #0282-0286 5 x (3 x 4.5) Halo1
N30 FeII #0287-0291 5 x (3 x 4.5) lamp off
N30 FeII #0292-0296 5 x (3 x 4.5) Halo1
N30 Br Gamma #0297-0301 5 x (3 x 4.5) lamp off
N30 Br Gamma #0302-0306 5 x (3 x 4.5) Halo1
N30 Ks #0307-0311 5 x (3 x 9) lamp off
N30 Ks #0312-0316 5 x (3 x 9) Halo2
N30 H #0317-0321 5 x (3 x 8.5) lamp off
N30 H #0322-0326 5 x (3 x 8.5) Halo2

-- %USERSIG{JochenHeidt - 2016-11-24}%
Topic revision: r1 - 24 Nov 2016, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback